how do u work out a pecentage of a number such as 5% of 80 or 25& of 100
i already know the easy ones like 50& and 25% but what are the other ones?

Found 2 solutions by Dymind, kiru_khandelwal:
Answer by Dymind(3)

to find the percentage of any number divide the percentage by 100 and multiply
it by the number. eg 5% of 80= 5/100*80=4

Answer by kiru_khandelwal(79)

Percentage means per 100
so 5% means = 5/100
so 5% 0f 80 will be (5/100)*80= 400/100 = 4
SImilarly u can find out
for example 15% of 60
so 15% of 60 = (15/100) of 60 = (15/100) * 60 = 900/100 = 9
Hope this helps....
Now let me explain the other way round
How much percentage of 60 is 9?
Let 9 be x% of 60
=> (x/100)*60 = 9
=> 60x = 900
=> x = 900/60 = 15
So if a person gets 9 marks out of 60....in percentage it will be (9/60)*100.